# FRABench and GenEval: Scaling Fine-Grained Aspect Evaluation across Tasks, Modalities

### v2

@@ Abstract
Evaluating the open-ended outputs of large language models (LLMs) has become a bottleneck as model capabilities, task diversity, and modality coverage rapidly expand. Existing “LLM-as-a-Judge” evaluators are typically narrow in a few tasks, aspects, or modalities, and easily suffer from low consistency. In this paper, we argue that explicit, fine-grained aspect specification is the key to both generalizability and objectivity in automated evaluation. To this end, we propose a hierarchical aspect taxonomy encompassing 112 distinct aspects that unifies evaluation across four representative settings — Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation. Building upon this taxonomy, we create FRABench, a benchmark comprising 60.4k pairwise samples with 325k evaluation labels obtained from a combination of human and LLM annotations. FRABench provides the first large-scale, multi-modal resource for training and meta-evaluating fine-grained LMM judges. Leveraging FRABench, we develop GenEval, a fine-grained evaluator generalizable across tasks and modalities. Experiments show that GenEval (i) attains high agreement with GPT-4o and expert annotators, (ii) transfers robustly to unseen tasks and modalities, and (iii) reveals systematic weaknesses of current LMMs on evaluation [^1] Our code can be found at https://github.com/ALEX-nlp/FRABench-and-GenEval.
1 Introduction
As large language models (LLMs) have shown amazing abilities in human-like question answering and creative generation [1, 2], assessing the quality of their free-form outputs has become increasingly challenging. Automated evaluators, a.k.a. “LLM-as-a-Judge” paradigm [3], have therefore received much research attention, with the purpose of providing absolute or relative scores sometimes accompanied by detailed feedback [4]. While existing evaluators [5, 6, 7] are typically tailored to limited tasks, aspects, or modalities, which sharply constrains their scope of application. Equally troubling, the credibility of these evaluators remains in doubt — empirical evidence shows that even human annotators suffer from a low agreement in evaluating the open-ended responses of LLMs as task complexity rises, and current automated evaluators inherit the same vulnerability.
In this paper, we target a fine-grained evaluator designed to generalize across different tasks and modalities (i.e., text and image). We highlight that aspect-level assessment is the core of generalizability. Because aspects like engagement, naturalness, or creativity are inherently interrelated [8], thus transferring similar semantics and evaluation standards for various tasks. Additionally, explicit aspect specification can also reduce subjectivity. Our observations indicate that the low agreement among human annotators is mainly caused by their implicitly varying weighting over aspects, rather than disagreement over the definition of any single aspect. For example, when judging story generation, some prioritize creativity while others emphasize logical consistency. By specifying aspects, one can substantially mitigate this variance source and obtain more objective evaluations.
Method Task Modalty Aspect Generalizable Data Source
@@ 1 Introduction
X-Eval [5] 27 Human
Themis [9] 50 Synthetic
ImageReward [10] 3 Human
VisionReward [11] 37 Human
LLaVA-Critic [12] Synthetic
GenEval 112 Synthetic & Human
Table 1: Comparison of our GenEval with recent evaluators. GenEval supports evaluating across four tasks and two modalities, incorporates multiple aspects, and demonstrates strong generalization ability. “–” indicates that the number of supported aspects is not explicitly specified.
To this end, we first conduct a comprehensive survey on four types of (multi-modal) tasks: Natural Language Generation (NLG), Image Understanding (IU), Image Generation (IG), and Interleaved Text-and-Image Generation (ITIG) to identify key evaluation aspects. We then manually organize, extend, and redefine 112 distinct aspects with hierarchical relations as a universal evaluation taxonomy—aspect tree. Note that although recent studies like Prometheus [13] and BIGGen-Bench [14] use auto-generated aspects for ultra-fine evaluation, we avoid this approach for two reasons. First, when the aspect becomes that narrowly tailored, its boundaries with direct evaluation become blurred, thereby reintroducing the very subjectivity we aim to suppress. Second, the universe of evaluation aspects is comparatively stable and, in practice, enumerable. Once the application scenario is specified, human or automated evaluators can simply traverse our aspect tree to select the relevant ones. This design dramatically lowers the cognitive load of defining evaluation standards and promotes consistency across tasks and modalities.
@@
Based on the aspect tree, we construct a large-scale dataset named Fine-grained Aspect Benchmark (FRABench) for both training and testing. We leverage the aspect tree to select relevant multiple aspects for each of the 60.4k pairwise responses across 28 sub-tasks. To obtain high-quality evaluation labels while maintaining cost efficiency, we collect available human annotations wherever possible and supplement them with GPT-4o-assisted completions. In total, this process yields 325k evaluation labels. Using FRABench, we develop GenEval, a generalizable evaluator. Table 1 shows the detailed comparison with existing evaluators. Experimental results demonstrate that GenEval exhibits strong generalization ability on both in-domain and out-of-domain tests, and achieves high consistency with GPT-4o and human annotations. Meanwhile, we also identify several limitations and deficiencies in current LMMs when evaluating IG tasks, which can serve as promising directions for future research.
In summary, our contributions are as follows: (1) We construct an evaluation aspect tree covering 112 aspects to ease fine-grained evaluation. (2) We develop FRABench, the first large-scale, multi-modal resource for training and meta-evaluating fine-grained LMM judges. (3) We create GenEval and conduct experiments to showcase its generalization capabilities, providing key insights.
2 Fine-grained Aspect Tree Construction
To ease fine-grained evaluation for both automated evaluators and humans, we construct an aspect tree that is universal across modalities and tasks. The construction pipeline involves two main steps: (1) Collecting and extending aspects. (2) Constructing aspect taxonomy.
2.1 Aspect Collection and Extension
@@
To fully leverage existing aspects, we first collect 28 sub-tasks under the four types of tasks (i.e., NLG, IU, IG, and ITIG). These sub-tasks span all six combinations of input types (text and text-with-image) and output types (text, image, and text-with-image), ensuring comprehensive coverage across multimodal tasks. Subsequently, we systematically collected literature related to each sub-task to gather relevant evaluation aspects and, where available, their definitions. For aspects without available definitions, we define them according to their practical meaning.
[…]
2.2 Aspect Taxonomy Construction
Figure 1: Universal Aspects: A subtree rooted at overall, driven by task independence.
Figure 2: Task-specific Aspects: A subtree rooted at overall, driven by task dependence.
In this section, we construct an aspect tree based on the aspects to establish a standardized evaluation taxonomy for multimodal evaluation tasks. Based on this universal taxonomy, both human and automated evaluators can easily traverse the aspect tree to customize evaluation settings for specific tasks or scenarios, enabling fine-grained and objective assessments.
@@ 2.2 Aspect Taxonomy Construction
We first designate the overall aspect as the root node and construct two subtrees, then categorize the remaining aspects into corresponding subtrees based on task independence and specificity:
@@
  • Universal Aspects (UAs): Aspects in this subtree exhibit task independence, as they primarily focus on assessing the quality of the model’s output. They are often modality-specific (for example, fluency for text or fidelity for image) and serve as fundamental quality aspects across tasks.
@@
  • Task-specific Aspects (TAs): Aspects in this subtree exhibit task dependence, as they primarily focus on task completion rather than output quality. They are typically closely tied to the task type. For example, engagingness for story generation or accuracy for mathematical reasoning.
For the hierarchical construction within the two subtrees, we first adopt existing aspect trees from prior studies [15, 16, 17, 18, 19, 20]. For aspects lacking predefined hierarchies, we adopt an iterative methodology where, in accordance with the definitions, the aspect is either (1) integrated into an existing tree (for example, the definition of alignment encompasses both object alignment and attribute alignment.) or (2) created as the root node of a new tree. This definition-driven approach is reasonable, as it ensures that the constructed hierarchy accurately reflects the semantic relationships among aspects. Finally, each tree’s root nodes serve as child nodes of the corresponding two subtrees. The final two subtrees are shown in Figure 2 and Figure 2.
Based on the aspect tree, the aspect-level evaluation process can be streamlined. Specifically, we first determine the correspondence between each subtree (e.g, the subtree rooted at readability, represented by readability) and its respective modality: readability for text, image quality for image, image coherence for multi-images, and text-image relationship for text-with-image, while bias and toxicity apply across all modalities. Once the evaluation task is determined, we can identify the corresponding subtree based on the output modality to select appropriate UAs. Additionally, task-specific aspects can be traversed from the TAs according to the objectives of the task.
3 FRABench Construction
To support fine-grained evaluation, we construct FRABench, a comprehensive dataset for training and evaluating evaluators on the aspect tree we proposed. Following prior studies [21, 22] that demonstrate pointwise scoring is more susceptible to contextual bias, we adopt a pairwise comparison evaluation method in FRABench. FRABench comprises 60.4K pairwise samples and 325k evaluation labels. Detailed statistics are presented in Figure 3 and Appendix B.4. Moreover, we rigorously partition the FRABench dataset into training, in-domain test, and out-of-domain test subsets to enable evaluation of both in-domain and out-of-domain generalization capabilities of evaluators.
@@ 3 FRABench Construction
Figure 3: Data statistics of the FRAbench. Besides using public datasets, we generated an equivalent amount of pairwise data. Detailed counts of evaluation labels for each sub-task are in Appendix B.4.
3.1 Dataset Collection
@@
We collect queries from the sub-tasks in the aspect tree we proposed and generate pairwise samples. Specifically, we obtain 29.3K response pairs from high-quality public datasets and generate the remaining 30.1K pairs using LMMs (model details provided in Appendix B.1). After obtaining the queries and response pairs, we assign aspects from our aspects tree(the aspect selection method is described in Section 2.2). As a result, each pairwise sample is annotated with an average of 8 UAs and 3 TAs. We then generate fine-grained evaluation labels for each response pair. Our evaluation labels have two sources: (1) Human annotations: We directly obtain human-annotated evaluation scores covering three IG-related aspects from the ImageRewardDB dataset [10], and supplement corresponding feedback using GPT-4o to form the evaluation labels. (2) GPT-4o annotations: For all other pairwise samples, due to the absence of human-annotated labels, we use GPT-4o to generate evaluation labels. All template details are provided in Appendix F. Moreover, to mitigate position bias commonly observed in evaluators [22], we balance the number of samples where response 1 is preferred over response 2 and vice versa. Finally, we generate 325k fine-grained evaluation labels.
3.2 Dataset Partitioning
To comprehensively train and evaluate the generality of GenEval, we split FRABench into training, in-domain, and out-of-domain test subsets. Specifically, we randomly select 18 sub-tasks from 4 tasks, covering 22 UAs and 35 TAs for the training set and in-domain test set (FRA-ID). Meanwhile, we select 10 unseen sub-tasks covering 28 seen UAs and 27 unseen TAs for the out-of-domain test set (FRA-OOD). Additionally, to assess GenEval’s generalization across UAs, we construct FRAUAs-OOD, which includes the same sub-tasks as the training set but different UAs across 4 tasks. Finally, the training set contains 255.4K samples, FRA-ID contains 45.2K samples, FRA-OOD contains 24.4K samples, and FRAUAs-OOD contains 5.3K samples.
@@ 3.2 Dataset Partitioning
The role of an evaluator is to assess model responses in place of humans. Therefore, the primary criterion for evaluating an evaluator’s effectiveness is their alignment with human judgments. Therefore, we construct corresponding evaluation datasets to measure the consistency between the evaluator’s assessments and human judgments. We extract partial samples from FRA-ID and FRA-OOD, and recruited three human annotators with master’s degrees or higher for annotation. Finally, we retain 6.9K in-domain samples and 6.0K out-of-domain samples with highly consistent human annotations, forming the final in-domain evaluation dataset (FRA-ID-H) and out-of-domain evaluation dataset (FRA-OOD-H), respectively. Details of the human annotation process are provided in Appendix E.
4 Evaluator Training
To train the GenEval, we use the training set from FRABench with Qwen2-VL-7B-Instruct as the backbone. As our primary focus is on fine-grained evaluation and the construction of the aspect tree and FRABench, we adopt the SFT for training. More advanced training techniques, such as RL, are left for future work. For training strategies, we explore the effectiveness of staged training versus unified training in the multi-task setting. Specifically, we split the training data by output modality (text, image, and text-image) and perform staged training in all six possible orders, producing six evaluators. The results indicate that phased training causes catastrophic forgetting. As a result, we adopt the unified training strategy. Details of the training setup and the results of the phased training experiments are presented in Appendix C.1 and Appendix D.2, respectively.
5 Experiments
To comprehensively evaluate GenEval, we consider three experimental settings: In-Domain Evaluation, Out-of-Domain Evaluation, and Evaluation as an Evaluator and Reward Model. Next, we present the experimental setup and corresponding results.
5.1 Experimental Setup
Benchmarks. (1) In-Domain Evaluation: we use FRA-ID and FRA-ID-H for assessment. (2) Out-of-Domain Evaluation: FRAUAs-ODD, FRA-OOD, and FRA-OOD-H are employed. We also include some public benchmarks: SummEval [23] and MANS [24] for NLG, WildVision [25] for IU, and GenAI-Bench [26] and Winoground [27] for IG. (3) Evaluation as Evaluator and Reward Model: MT-Bench [28] and MLLM-as-a-Judge [29] are used to evaluate models as an evaluator. VLRewardBench [30] is used for reward model evaluation. Detailed descriptions of each benchmark are provided in Appendix C.2.
Baselines. we use following models as baselines: GPT-4o, Claude-3.5, Qwen2-VL-72B-Instruct (Qwen2VL-72B), Qwen2-VL-7B-Instruct (Qwen2VL-7B), Themis-8B [9] (Themis), LLaVA-Critic-7B [12] (LLaVA-Critic), ImageReward [10] and VisionReward [11]. Notably, the paradigm adopted by ImageReward and VisionReward for fine-grained evaluation differs from ours. Therefore, they are used solely as baselines in public benchmark evaluations.
5.2 In-Domain Evaluation
For in-domain evaluation, we assess performance using the FRA-ID and FRA-ID-H. Figure 4 presents radar charts that illustrate the overall degree of alignment [^3] We use aspect abbreviations here; the full names can be found in Appendix A.. Detailed scores for each aspect are reported in Appendix D.1. The accompanying Table 2 reports average accuracies for each task. Additionally, we also sample three types of aspects for aspect-level analysis. As shown in Figure 5, we select aspects from four test sets based on GenEval’s performance: (1) Weak aspects, where the best model surpasses GenEval by over 20%. (2) Universal poor aspects, where scores from all evaluators are very low. (3) Strong aspects, where GenEval excels. Only two weak aspects appear across all test sets, and both are included. For the other categories, we choose representative subsets.
The results demonstrate that GenEval consistently achieves higher alignment with both GPT-4o and human annotators compared to the baselines (see shaded area in radar chart). Specifically, excluding GPT-4o, GenEval surpasses all baselines on NLG, IU, and ITIG tasks within FRA-ID and FRA-ID-H, achieving overall average accuracies of 82.4% and 72.1%, respectively. On NLG, GenEval outperforms Themis, while our base model Qwen2Vl-7B performs much worse, suggesting that with appropriate training, LMM-based evaluators can also effectively assess NLG tasks. Moreover, GenEval shows strong performance on the ITIG of FRA-ID-H, where its alignment with human annotators exceeds that of GPT-4o. This improvement can be attributed to GenEval’s generalization ability and cross-task transfer performance. For aspect-level comparison, GenEval outperforms GPT-4o on some aspects, with two illustrative examples shown in Figure 5a.
Method Seen UAs for Seen Tasks Seen TAs for Seen Tasks
GPT-4o (FRA-ID) Human (FRA-ID-H) GPT-4o (FRA-ID) Human (FRA-ID-H)
@@ 5.2 In-Domain Evaluation
GPT-4o 94.0 91.7 85.5 90.9 76.1 79.1 65.0 66.1 92.5 90.0 76.6 89.4 79.5 74.8 58.8 73.6
Themis 60.4 - - - 60.3 - - - 47.5 - - - 53.4 - - -
LLaVA-Critic - 47.5 - - - 39.0 - - - 68.0 - - - 56.0 - -
We also observe special phenomena. As shown in Table 2, IG has the lowest accuracy among the four tasks across all baselines. In aspect analysis, action alignment and object alignment in IG under FRA-ID-H receive consistently low scores across all baselines (Figure 5b). We will discuss the reasons in Section 5.3.1. Additionally, for the harmfulness of IG in FRA-IN (Figure 5c), GenEval demonstrates relatively poor performance. Our analysis indicates that this issue mainly arises because GenEval classifies shadowy or gloomy images as harmful due to its heightened sensitivity to elements that could evoke psychological discomfort. Relevant cases are presented in Appendix G.
Figure 6: Comparison on FRA-OOD and FRA-OOD-H. Black sectors show unseen TAs (for task generalization), while the rest show seen UAs for unseen tasks (for task generalization).
Method Seen UAs for Unseen Tasks Unseen TAs for Unseen Tasks
GPT4o (FRA-OOD) Human (FRA-OOD-H) GPT4o (FRA-OOD) Human (FRA-OOD-H)
NLG IU IG ITIG NLG IU IG ITIG NLG IU IG ITIG NLG IU IG ITIG
GPT-4o 90.8 94.7 89.9 88.2 84.0 82.1 72.3 93.1 87.3 89.7 85.9 91.0 83.2 82.1 74.2 93.1

### v3

# UFEval: Unified Fine-grained Evaluation with Task and Aspect Generalization

@@ Abstract
Evaluating the open-ended outputs of Large Multimodal Models has become a bottleneck as model capabilities, task diversity, and modality rapidly expand. Existing "LLM-as-a-Judge" evaluators are typically narrow in specific tasks and aspects. In this paper, we argue that, on one hand, based on the interconnected nature of aspects, learning specific aspects can generalize to unseen aspects; on the other hand, jointly learning to assess multiple visual aspects and tasks may foster a synergistic effect. To this end, we propose UFEval, the first unified fine-grained evaluator with task and aspect generalization for four evaluation tasks — Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation. Specifically, (1) We first construct a hierarchical aspect taxonomy encompassing 112 distinct aspects across the aforementioned four tasks. (2) Then, building upon this taxonomy, we create FRABench, a fine-grained evaluation dataset comprising 60.4k pairwise samples with 325k evaluation labels obtained from a combination of human and GPT-4o annotations. FRABench provides a large-scale, multi-modal, and aspect-level resource for training and testing evaluators. (3) Finally, leveraging FRABench, we develop UFEval, a unified fine-grained evaluator. Experiments show that learning on specific aspects enables UFEval to generalize to unseen aspects, and joint learning to assess diverse tasks and aspects can lead to substantial mutual benefits.
Codehttps://github.com/ALEX-nlp/UFEval
Datasetshttps://huggingface.co/datasets/SPUH/FRABench
Introduction
As Large Multimodal Models have shown amazing abilities in human-like question answering (Grattafiori et al. 2024; Wang et al. 2024a), assessing the quality of their free-form outputs has become increasingly challenging. Automated evaluators, a.k.a. “LLM-as-a-Judge” paradigm (Saha et al. 2025) have therefore received much research attention (Li et al. 2024b). Despite their progress, we posit two concerns: (1) Current evaluators (Liu et al. 2023b; Ke et al. 2023) are typically tailored to specific aspects, which limits their adaptability to unseen aspects. (2) They are also limited to specific tasks and modalities, which sharply constrains their scope of application. Table 1 shows the detailed comparison with existing evaluators. We intuitively argue that, on one hand, evaluation aspects are inherently interconnected (Fu et al. 2023). Specifically, aspects such as engagement, naturalness, and creativity are closely linked. Thus, similar semantics and evaluation standards can be transferred across diverse tasks. On the other hand, jointly learning to assess multiple visual aspects and tasks may foster synergistic effects (Wang et al. 2025). For instance, learning object alignment in image captioning improves character consistency evaluation in multi-image scenarios through better object recognition. Furthermore, progress in image understanding tasks can enhance the evaluation of image generation results by enabling more accurate judgments of content quality and contextual appropriateness. This cross-aspect and cross-task synergy enables a more robust evaluation for diverse outputs. It inspires the development of a unified, fine-grained evaluator to achieve better aspect of generalization and evaluation performance.
Method Task Modalty
@@ Introduction
Auto-J
X-Eval
Prometheus 2
Themis
LLaVA-Critic
ImageReward
VisionReward
UFEval (Ours)
Table 1: Comparison of our UFEval with recent representative evaluators, including Auto-J (Li et al. 2023b), X-Eval (Liu et al. 2023b), Prometheus 2 (Kim et al. 2024a), Themis (Hu et al. 2024b), LLaVA-Critic (Xiong et al. 2024), ImageReward (Xu et al. 2023a), and VisionReward (Xu et al. 2024). NLG, IU, IG, and ITIG represent Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation, respectively.
@@
To this end, we propose UFEval, the first unified fine-grained evaluator with task and aspect generalization for four evaluation tasks (i.e., NLG, IU, IG, and ITIG) across 28 sub-tasks. However, training such a unified evaluator requires large-scale, multi-modal, aspect-level evaluation datasets, which are currently unavailable. Therefore, we develop the Fine-grained Aspect Benchmark (FRABench) to address this gap. Specifically, we first conduct a survey on the four tasks to identify key evaluation aspects. We then manually organize, extend, and redefine 112 distinct aspects with hierarchical relations as a universal evaluation taxonomy—aspect tree. Leveraging the aspect tree, we construct FRABench to ensure comprehensive aspect coverage. We leverage the aspect tree to select relevant multiple aspects for each of the 60.4k pairwise responses. To obtain high-quality evaluation labels while maintaining cost efficiency, we collect available human annotations wherever possible and supplement them with GPT-4o-assisted completions. In total, this process yields 325k evaluation labels. Using FRABench, we develop UFEval, a unified and fine-grained evaluator.
Our experiments demonstrate that UFEval exhibits excellent evaluation quality and strong task and aspect generalization capabilities. This is attributed to learning multiple aspects and tasks jointly, which yields significant mutual enhancement, as well as to fully leveraging the interconnections between learning and unlearning aspects. By conducting ablation experiments across diverse aspects and task baselines, we observe a gradual improvement in evaluation quality, thereby validating our hypothesis.
In summary, our contributions are as follows: (1) We construct FRABench, the large-scale multi-modal aspect-level evaluation dataset. (2) We develop UFEval, the first unified fine-grained evaluator for multiple tasks assessment with task and aspect generalization. (3) Our experiments show that inter-aspect correlations enable generalizable capability, and learning to assess multiple tasks and aspects jointly leads to a synergistic improvement in evaluation performance.
Related Work
Coarse-grained Specific-task Evaluation
Previously, coarse-grained evaluators, which produce overall judgments based on one or several aspects, have been widely explored across various evaluation tasks (Zhu, Wang, and Wang 2023; Xu et al. 2023b; Ye et al. 2023; Liu et al. 2024b; Wang et al. 2023a; Jiang et al. 2023). For instance, in NLG, Wang et al. (Wang et al. 2023b) first introduce PandaLM, a fine-tuned LLM designed to evaluate pairwise texts based on several aspect. Similarly, Li et al. (Li et al. 2023b) propose Auto-J, an evaluator capable of handling a broader range of tasks and aspects, supporting both pointwise and pairwise evaluation settings. In image generation evaluation, Xu et al. (Xu et al. 2023a) develop ImageReward, an evaluator trained on a large-scale human annotation dataset. Using alignment and fidelity as reference dimensions, ImageReward provides an overall assessment. Coarse-grained evaluations, while effective overall, lack the granularity to identify specific model weaknesses and often introduce aspect bias.
Fine-grained Specific-task Evaluation
To address the limitations of coarse-grained evaluation, recent studies have shifted toward fine-grained evaluation (Ye et al. 2023; Kim et al. 2024b; Ke et al. 2023; Kim et al. 2023; Li et al. 2023d; Ying et al. 2024; Bai et al. 2023), where LMs are fine-tuned on multi-aspect datasets to produce aspect-specific judgments. For example, in NLG, Hu et al.(Hu et al. 2024b) propose Themis, an LLM trained on the GPT-4 annotated NLG-Eval corpus. In IU, LLaVA-Critic(Xiong et al. 2024) is the first fine-grained evaluator integrating diverse criteria, showing strong correlation with GPT-4o. For IG, Xu et al.(Xu et al. 2024) introduce VisionReward, a VQA-based evaluator that assesses image quality across fine-grained aspects. While these methods improve over coarse-grained evaluation by targeting specific aspects, they struggle with scalability across different tasks and aspects. Liu et al.(Liu et al. 2023b) further investigate aspect generalization through X-Eval, a two-stage learning framework that incorporates auxiliary aspects and demonstrates generalization in NLG. However, its reliance on predefined reference aspects and lack of open-source implementation restricts its usability and reproducibility.
Methodology
To develop UFEval, we need to construct a large-scale, multi-modal, and aspect-level evaluation dataset, which is currently unavailable. The construction pipeline involves two main steps: (1) Evaluation Aspect Construction, and (2) Fine-grained Evaluation Dataset Construction. The following sections detail these two steps.
Evaluation Aspect Construction
Aspect Collection and Extension.
@@
To fully leverage existing aspects, we first collect 28 sub-tasks under the four types of tasks (i.e., NLG, IU, IG, and ITIG). These sub-tasks span all six combinations of input types (text and text-with-image) and output types (text, image, and text-with-image), ensuring comprehensive coverage across multimodal tasks. Subsequently, we collect literature related to each sub-task to gather relevant aspects and, where available, their definitions. Moreover, we define aspects without available definitions according to their practical meaning.
Beyond collecting existing aspects, we extend aspects for the ITIG, which lacks sufficient aspects. To this end, we apply a cross-task transfer by identifying analogous sub-tasks from other categories and adapting their aspects. For example, both story generation (NLG) and visual story completion (ITIG) involve narrative creation, enabling aspects like engagingness to be adapted  [^1] Aspects sharing the same term but differing in definitions are sequentially marked with superscripts: $\dagger$} and * .. Ultimately, we obtain 112 different aspects, with detailed information and sources listed in Appendix A.
Figure 1: An illustration of the evaluation pipeline. The pipeline consists of two steps: (1) Aspect Selection: First, appropriate aspects are selected from the TAs and UAs Trees based on task property and output modality. As illustrated, the question focuses on “cat image next to the orange dog picture”, while the image shows no cat adjacent to the orange dog. This discrepancy enables evaluation of the model’s ability to avoid hallucinations, specifically its performance on Context Inconsistency. Second, given the text output modality, aspects are selected from the Text Branch in the UAs Tree. (2) Evaluating: UFEval generates comprehensive feedback and relative scores based on the input content and selected aspects.
Aspect Taxonomy Construction.
@@
To facilitate aspect selection within the dataset construction and evaluation pipeline, we organize the collected aspects into an aspect tree serving as a standardized taxonomy. We first use the “overall” aspect as the root node of the aspect tree, and then divide the remaining aspects into two subtrees based on their task independence and task specificity:
@@
  • Universal Aspects (UAs): These aspects exhibit task independence, as they primarily focus on assessing the quality of the model’s output. They are often modality-specific (for example, fluency for text or fidelity for image).
@@
  • Task-specific Aspects (TAs): These aspects exhibit task dependence, as they primarily focus on task property rather than output quality. For example, engagingness for story generation or accuracy for mathematical reasoning.
@@
Figure 2: Data statistics of the FRAbench. FRABench comprises the largest number of evaluation aspects among existing aspect-level evaluation datasets. Detailed counts of evaluation labels for each sub-task are in Appendix B.
For hierarchical construction within the two subtrees, we directly adopt existing aspect tree structures from previous studies (Hu et al. 2024a; Ghosh, Hajishirzi, and Schmidt 2023; Huang et al. 2023; DeAlcala et al. 2023; Zeng et al. 2023; Huang et al. 2025). For aspects lacking predefined hierarchies, we determine their positions within the tree based on their definitions—either assigning them to locations within existing structures or establishing them as leaf nodes of root node. Based on the aspect tree, the fine-grained evaluation process can be streamlined, as illustrated in Figure 1.
Fine-grained Evaluation Dataset Construction
Based on the aspect tree, we can construct fine-grained evaluation datasets for training and testing UFEval. Specifically, we construct FRABench—a large-scale, multi-modal, and aspect-level evaluation dataset. Following prior studies (Kim et al. 2024a; Ye et al. 2024) that demonstrate pointwise scoring is more susceptible to contextual bias and available for reward model training, we adopt a pairwise comparison evaluation method in FRABench.
FRABench Construction.
@@
We first collect questions from the datasets corresponding to the 28 sub-tasks we gathered, and generate pairwise samples. Specifically, we first obtain 29.3K response pairs from public datasets and generate the remaining 30.1K pairs using LMMs (model details provided in Appendix B). After obtaining the queries and response pairs, we assign aspects from our aspects tree. As a result, each pairwise sample is annotated with an average of 8 UAs and 3 TAs (Aspect selection information in Appendix B). We then generate evaluation labels via two approaches: (1) Human annotations: directly using three-aspect human-annotated scores from ImageRewardDB (Xu et al. 2023a), supplemented with feedback generated by GPT-4o; (2) GPT-4o annotations: due to the lack of human annotations for most aspects, we use GPT-4o to generate evaluation labels for all other pairwise samples. All templates are provided in Appendix F. Moreover, to mitigate position bias (Ye et al. 2024), we balance the number of samples where response 1 is preferred over response 2 and vice versa by reversing the response positions in half of the surplus samples from the majority class and re-annotating them accordingly. Finally, we generate 325k fine-grained evaluation labels.
Dataset Partitioning and Evaluator Training.
Due to the lack of a large-scale aspect-level meta-benchmark to verify UFEval’s aspect generalizable capability and support our arguments, we can only divide FRABench into training, In-Domain test set (FRA-ID), and Out-of-Domain test set (FRA-OOD). The training and FRA-ID consist of 18 randomly selected sub-tasks from four tasks, covering 22 UAs and 35 TAs, while FRA-OOD contains 10 unseen sub-tasks with 28 seen UAs and 27 unseen TAs. Finally, the training set contains 255.4K samples, FRA-ID contains 45.2K samples, and FRA-OOD contains 24.4K samples. Detailed statistics are presented in Figure 2 and Appendix B.
@@
The role of an evaluator is to assess model responses in place of humans. Therefore, the primary criterion for evaluating an evaluator’s effectiveness is their alignment with human judgments. To this end, we construct corresponding evaluation datasets to measure the consistency. We extracted partial samples from FRA-ID and FRA-OOD, and recruited three humans with master’s degrees for annotation. Finally, we retain 6.9K in-domain evaluation samples (FRA-ID-H) and 6.0K out-of-domain evaluation samples (FRA-OOD-H), respectively. Details of the human annotation process and annotation consistency are provided in Appendix E.
When training UFEval, we use SFT to fine-tune Qwen2-VL-7B-Instruct on the training set. The detailed training configurations and methods are presented in Appendix C.
Experiments
To assess UFEval’s generalizability stemming from inter-aspect correlations, we conduct evaluations on previously unseen tasks and aspects, termed Out-of-Domain Evaluation. We then evaluate UFEval as LLM-as-a-Judge using public meta-benchmarks. Finally, we explore the effectiveness of Multi-aspect and Multi-task Assessment Learning for evaluators. In the following sections, we present our experimental setup and corresponding results (In-domain and other experimental results are in Appendix D).
Figure 3: Comparison of baselines on the FRA-OOD and FRA-OOD-H. Black sectors show unseen TAs (for aspect generalization), while the rest show seen UAs for unseen tasks (for task generalization).
Method Tasks Generalization Evaluation Aspects Generalization Evaluation
FRA-OOD (GPT4o) FRA-OOD-H (Human) FRA-OOD (GPT4o) FRA-OOD-H (Human)
@@
GPT-4o - - - - 84.0 82.1 72.3 93.1 - - - - 83.2 82.1 74.2 93.1